# Edge Importance in Complex Networks

### v1

@@ 4.1 Medium-sized networks
Table 5: Example 4.2. The first $10$ highway segments that should be widened/narrowed, in order to increase/decrease the network communication in Autobahn (in the first column) and the first $10$ highway segments that should be built in order to increase the network communication in Autobahn (in the second column), according to the edge importance vector ${\mathcal{E}}_{L_{f}}$ and the virtual edge importance vector ^${\widehat{\mathcal{E}}}_{L_{f}}$, respectively.
${\mathcal{E}}_{\rho}$ $e(v_{i}\leftrightarrow v_{j})$${\widehat{\mathcal{E}}}_{\rho}$ $e(v_{i}\leftrightarrow v_{j})$
[…]
Table 6: Example 4.2. The first $10$ highway segments that should be widened/narrowed, in order to increase/decrease the network communication in Autobahn (in the first column) and the first $10$ highway segments that should be built in order to increase the network communication in Autobahn (in the second column), according to the Perron edge importance vector ${\mathcal{E}}_{\rho}$ and the Perron virtual edge importance vector ^${\widehat{\mathcal{E}}}_{\rho}$, respectively.
EXAMPLE 4.3.
Consider the directed weighted graph $\mathcal{G}$ that represents the network C.elegans available at [28], i.e., the metabolic network of the Caenorhabditis elegans worm. The network contains $n=306$ vertices that represent neurons and $m=2345$ edges. Two neurons are connected if at least one synapse or gap junction exists between them and the associated edge-weight is the number of synapses and gap junctions. The network is disconnected.
The total communicability in (2) is $T_{\mathcal{G}}=3.3401\cdot 10^{6}$. The gradient $\nabla T_{\mathcal{G}}$ in (3) has been computed by evaluating the matrix $L_{f}(\mathbf{A}^{T},uv^{T})$ in (8) using (9). The total transmission is $\|\nabla T_{\mathcal{G}}\|_{2}=7.9032\cdot 10^{6}$. The gradient $\nabla T_{\mathcal{G}}$ has been approximated by evaluating $L_{f}(\mathbf{A}^{T},uv^{T})$ using (10) with $h=\frac{2}{n}\cdot 10^{-4}=6.5359\cdot 10^{-7}$, obtaining $\widetilde{\nabla T}_{\mathcal{G}}$. The resulting total transmission is $\|\widetilde{\nabla T}_{\mathcal{G}}\|_{2}=7.9032\cdot 10^{6}$, having
@@
As for both the edge importance vector ${\mathcal{E}}_{L_{f}}$ and the virtual edge importance vector ^$\widehat{\mathcal{E}}_{L_{f}}$, one obtains the same results, displayed in Table 7, regardless of whether $\nabla T_{\mathcal{G}}$ or $\widetilde{\nabla T}_{\mathcal{G}}$ is used.
Removing the two edges in bold face in the second column of Table 7, associated with the smallest entries of the edge importance results in the network ${\mathcal{G}}_{1}$ one obtains
@@
Setting to zero the weights associated with the two edges in bold face in the fourth column of Table 7) results in the network ${\mathcal{G}}_{2}$, for which one has
@@
Finally, setting to one the (vanishing) entries of $A$ associated with the two virtual edges in bold face in the sixth column of Table 7) returns the network ${\mathcal{G}}_{4}$, with
@@
Turning to the network modifications based on Perron root sensitivity in C.elegans, one has $P_{\mathcal{G}}=9.1975\cdot 10^{5}$. The graph ${\mathcal{G}}$ is reducible, some entries of the Perron vector $x$ are vanishing. In particular, removing the edges in bold face in the second column of Table 8, which are associated with two vanishing entries of the Perron edge importance vector, results in the network ${\mathcal{G}}_{5}$, for which one has
@@
So the simplification is less satisfactory than that obtained in ${\mathcal{G}}_{1}$. Elimination of the edges in bold face in the fourth column of Table 8) returns the network ${\mathcal{G}}_{6}$, for which one has
@@
The decrease is less than that in ${\mathcal{G}}_{2}$. Conversely, increasing by one the weights associated with the same edges results in the network ${\mathcal{G}}_{7}$, for which one has
@@
The increase is less than that in ${\mathcal{G}}_{3}$. Finally, setting to one the (vanishing) entries of $A$ associated with the two virtual edges in bold face in the sixth column of Table 8) results in the network ${\mathcal{G}}_{8}$, having
@@
Therefore the result is more satisfactory than that obtained in the network ${\mathcal{G}}_{4}$. Notice that $\|\mathbf{W}|_{{\mathcal{A}}}\|_{F}=1.2817\cdot 10^{-1}$.
@@
Table 7: Example 4.3. The smallest entries of the edge importance vector and the relevant junctions that could be removed without changing the network communication in C.elegans significantly (displayed in the first two columns). The largest entries of the edge importance vector and the relevant junctions to increase/decrease in order to increase/decrease the network communication (in the third and fourth columns). The largest entries of the virtual edge importance vector and the relevant junctions to add in order to increase the communication in C.elegans (in the fifth and sixth columns).
@@
Table 8: Example 4.3. The smallest entries of the Perron edge importance vector and the relevant junctions that could be removed without changing the communication in C.elegans significantly (displayed in the first two columns). The largest entries of the Perron edge importance vector and the relevant junctions to increase/decrease in order to increase/decrease the network communication in C.elegans (in the third and fourth columns). The largest entries of the virtual Perron edge importance vector and the relevant junctions to add in order to increase the network communication (displayed in the fifth and sixth columns).
@@ 4.2 Large networks
EXAMPLE 4.4.
Consider the unweighted undirected graph $\mathcal{G}$ that represents the continental US road network Usroads-48. The graph $\mathcal{G}$, which is available at [28], has $n=126146$ vertices, which represent intersections and road endpoints. The $m=161950$ edges represent roads which connect the intersections and endpoints. We analyze the network Usroads-48 with the tools discussed in Subsections 2.2 and 3.2 for large networks.
We would like to determine approximations of the smallest and largest elements of $L_{f}(A,ee^{T})|_{\mathcal{L}}$ that correspond to edges that should be removed to simplify the network or whose edge-weight should be modified to increase or decrease the total communicability. Moreover, we would like to determine approximations of the smallest and largest elements of $L_{f}(A,ee^{T})|_{\widehat{\mathcal{L}}}$ that correspond to edges that should be added to increase total communicability. We first carry out $5$ steps of the symmetric Lanczos process, computing $\mathbf{V}_{5}$ and $\mathbf{T}_{5}$, and make use of the latter to construct both $\mathbf{X}_{5}$ in (17), thanks to (18), and
@@
with $h=\frac{2}{n}\cdot 10^{-4}$. Then, proceeding as discussed in Subsection 2.2, both the approaches give the same estimates. The smallest element of the approximate $L_{f}(A,ee^{T})|_{{\mathcal{L}}}$ is $9.7462\cdot 10^{-7}$ and is associated with edge $e(v_{123259}\leftrightarrow v_{123258})$ while its largest element is $1.5008\cdot 10^{1}$ and is associated with edge $e(v_{19694}\leftrightarrow v_{19186})$; the smallest element of the approximate $L_{f}(A,ee^{T})|_{\widehat{\mathcal{L}}}$ is $2.2047\cdot 10^{-9}$ and is associated with edge $e(v_{25416}\leftrightarrow v_{11651})$ while its largest element is $1.9380\cdot 10^{1}$ and is associated with edge $e(v_{58080}\leftrightarrow v_{1})$.
Turning to the the structured Perron communicability, one has $P_{\mathcal{G}}^{{\mathcal{A}}}=1.9138\cdot 10^{2}$. The smallest entries of the Perron edge importance vector ${\mathcal{E}}_{\rho}$ and the relevant edges are displayed in Table 9 in the first and second columns, while the largest entries of ${\mathcal{E}}_{\rho}$ and the relevant edges are shown in the third and fourth columns. In order to increase the network communication one should add edge $e(v_{44182}\leftrightarrow v_{44035})$, which is associated with the largest entry of the $\widehat{\mathcal{L}}$-analogue of the Wilkinson perturbation.
@@
Table 9: Example 4.4. The smallest entries of the Perron edge importance vector and the edges associated with the roads that could be removed in order to reduce the complexity of the network without changing the communication in Usroads-48 significantly (displayed in the first two columns). The largest entries of the Perron edge importance vector and the edges associated with the first ten roads that should be widened/narrowed in order to increase/decrease the network communication the most (in the third and fourth columns).
Removing the two edges in bold face in the second column of Table 9, associated with the smallest entries of the edge importance results in the network ${\mathcal{G}}_{1}$, for which one has
$P_{{\mathcal{G}}_{1}}^{{\mathcal{A}}}=1.9138\cdot 10^{2};\;\;\;\;\frac{P_{\mathcal{G}}^{{\mathcal{A}}}-P_{{\mathcal{G}}_{1}}^{{\mathcal{A}}}}{P_{\mathcal{G}}^{{\mathcal{A}}}}=1.4851\cdot 10^{-16}.$
Setting to zero the weights associated with the two edges in bold face in the fourth column of Table 9) results in the network ${\mathcal{G}}_{2}$, having
$P_{{\mathcal{G}}_{2}}^{{\mathcal{A}}}=1.7487\cdot 10^{2};\;\;\;\;\frac{P_{\mathcal{G}}^{{\mathcal{A}}}-P_{{\mathcal{G}}_{2}}^{{\mathcal{A}}}}{P_{\mathcal{G}}^{{\mathcal{A}}}}=8.6228\cdot 10^{-2},$
$P_{{\mathcal{G}}_{3}}^{{\mathcal{A}}}=3.6445\cdot 10^{2};\;\;\;\;\frac{P_{{\mathcal{G}}_{3}}^{{\mathcal{A}}}-P_{\mathcal{G}}^{{\mathcal{A}}}}{P_{\mathcal{G}}^{{\mathcal{A}}}}=9.0435\cdot 10^{-1}.$
@@
$P_{{\mathcal{G}}_{4}}^{{\mathcal{A}}}=2.4540\cdot 10^{2};\;\;\;\;\frac{P_{{\mathcal{G}}_{4}}^{{\mathcal{A}}}-P_{\mathcal{G}}^{{\mathcal{A}}}}{P_{\mathcal{G}}^{{\mathcal{A}}}}=2.8228\cdot 10^{-1}.$
@@ 5 Conclusion and comments on related work
The identification of important and unimportant edges is a fundamental problem in network analysis. Several techniques for this purpose have been described in the literature; see, e.g., [7, 8, 5, 10, 21, 22]. In [4] the authors propose a method that uses the gradient of the total communicability, and Schweitzer [27] recently described how the computational effort required by this method can be reduced. Section 2 of this paper reviews this method and discusses computational aspects when this technique is applied to small and medium-sized networks, as well as to large-scale networks. In particular, further ways to speed up the computations when the method is applied to large-scale networks are described.
Another approach to identify important and unimportant edges is to determine edge weights whose modification yields a relatively large change in the Perron root of the adjacency matrix. This is described in [10, 21]. The computations required are quite straightforward and the method discussed in the latter reference is easy to implement also for large-scale problems. We therefore are interest in whether modifications of the weights of the edges identified by this technique give a relatively large change in the total communicability. Section 3 reviews the method described in [21] and extends it to include edge removal. Computed examples reported in Section 4 show that, indeed, modifications of edge weights identified by the technique discussed in [21] yield relatively large changes in the total communicability.
Acknowledgments
Research by SN was partially supported by a grant from SAPIENZA Università di Roma, by INdAM-GNCS, and by the PRIN 2022 research project “Inverse Problems in the Imaging Sciences (IPIS)”, grant n. 2022ANC8HL.

### v2

@@ 4.2 Medium-sized networks
Table 5: Example 3. The first $10$ highway segments that should be widened/narrowed in order to increase/decrease the network communicability in Autobahn according to the edge importance vector ${\mathcal{E}}_{L_{f}}$ (in the first column) and according to the Perron edge importance vector ${\mathcal{E}}_{\rho}$ (in the second column), respectively.
Example 4
Consider the directed weighted graph $\mathcal{G}$ that represents the network C.elegans available at celegans306_usroads48 , i.e., the metabolic network of the Caenorhabditis elegans worm. The network contains $n=306$ vertices that represent neurons and $m=2345$ edges. Two neurons are connected if at least one synapse exists between them and the associated edge-weight is the number of synapses. The network is disconnected.
The total communicability (2) is $T_{\mathcal{G}}=3.3401\cdot 10^{6}$. The gradient $\nabla T_{\mathcal{G}}$ (3) has been computed by evaluating the matrix $L_{f}(\mathbf{A}^{T},uv^{T})$ in (8) using (9). The total transmission is $\|\nabla T_{\mathcal{G}}\|_{2}=7.9032\cdot 10^{6}$. The gradient $\nabla T_{\mathcal{G}}$ has been approximated by evaluating $L_{f}(\mathbf{A}^{T},uv^{T})$ using (10) with $h=\frac{2}{n}\cdot 10^{-4}=6.5359\cdot 10^{-7}$ to obtain $\widetilde{\nabla T}_{\mathcal{G}}$. The resulting total transmission is $\|\widetilde{\nabla T}_{\mathcal{G}}\|_{2}=7.9032\cdot 10^{6}$, having
@@
As for both the edge importance vector ${\mathcal{E}}_{L_{f}}$ and the virtual edge importance vector ^$\widehat{\mathcal{E}}_{L_{f}}$, one obtains the same results, displayed in Table 6, regardless of whether $\nabla T_{\mathcal{G}}$ or $\widetilde{\nabla T}_{\mathcal{G}}$ is used. We remark that evaluating $L_{f}(\mathbf{A}^{T},ee^{T})$ by (9) required the average time $t_{1}\approx 1.3\cdot 10^{-1}$ in $10^{4}$ tests, while evaluating $L_{f}(\mathbf{A}^{T},ee^{T})$ using (10) required the average time $t_{2}\approx 1.7\cdot 10^{-2}$ in the same tests. The relative average time saving is
$\frac{t_{1}-t_{2}}{t_{1}}\approx 8.6\cdot 10^{-1}.$
Removing the two edges in bold face in the second column of Table 6, associated with the smallest entries of the edge importance, results in the network ${\mathcal{G}}_{1}$ for which
@@
Setting to zero the weights associated with the two edges in bold face in the fourth column of Table 6 results in the network ${\mathcal{G}}_{2}$, for which one has
@@
while increasing by one the weights associated with the same edges results in the network ${\mathcal{G}}_{3}$ with
@@
Finally, setting to one the (vanishing) entries of $A$ associated with the two virtual edges in bold face in the sixth column of Table 6 yields the network ${\mathcal{G}}_{4}$, for which
@@
Turning to network modifications based on Perron root sensitivity in C.elegans, one has $P_{\mathcal{G}}=9.1975\cdot 10^{5}$. The graph ${\mathcal{G}}$ is reducible, some entries of the Perron vector $x$ are vanishing. In particular, removing the edges in bold face in the second column of Table 7, which are associated with two vanishing entries of the Perron edge importance vector, results in the network ${\mathcal{G}}_{5}$, for which one has
@@
Hence, this simplification is less satisfactory than ${\mathcal{G}}_{1}$. Elimination of the edges in bold face in the fourth column of Table 7 gives the network ${\mathcal{G}}_{6}$, for which one has
@@
The decrease is less than that for ${\mathcal{G}}_{2}$. Conversely, increasing by one the weights associated with the same edges results in the network ${\mathcal{G}}_{7}$, for which one has
@@
The increase is less than that for ${\mathcal{G}}_{3}$. Finally, setting to one the (vanishing) entries of $A$ associated with the two virtual edges in bold face in the sixth column of Table 7 results in the network ${\mathcal{G}}_{8}$, having
@@
Therefore the result is more satisfactory than that obtained for the network ${\mathcal{G}}_{4}$. Notice that $\|\mathbf{W}|_{{\mathcal{A}}}\|_{F}=1.2817\cdot 10^{-1}$.
@@
Table 6: Example 4. The smallest entries of the edge importance vector and the relevant junctions that could be removed without changing the network communicability in C.elegans significantly (displayed in the first two columns). The largest entries of the edge importance vector and the relevant junctions to increase/decrease in order to increase/decrease the network communicability (in the third and fourth columns). The largest entries of the virtual edge importance vector and the relevant junctions to add in order to increase the communicability in C.elegans (in the fifth and sixth columns).
@@
Table 7: Example 4. The smallest entries of the Perron edge importance vector and the relevant junctions that could be removed without changing the communicability in C.elegans significantly (displayed in the first two columns). The largest entries of the Perron edge importance vector and the relevant junctions to increase/decrease in order to increase/decrease the network communicability in C.elegans (in the third and fourth columns). The largest entries of the virtual Perron edge importance vector and the relevant junctions to add in order to increase the network communicability (displayed in the fifth and sixth columns).
@@ 4.3 Large networks
Example 5
Consider the unweighted undirected graph $\mathcal{G}$ that represents the continental US road network Usroads-48. The graph $\mathcal{G}$, which is available at celegans306_usroads48 , has $n=126146$ vertices, which represent intersections and road endpoints. The $m=161950$ edges represent roads that connect the intersections and endpoints. We analyze the network Usroads-48 with the tools discussed in Subsections 2.2 and 3.2 for large networks.
We would like to determine approximations of the smallest and largest elements of $L_{f}(A,ee^{T})|_{\mathcal{L}}$ that correspond to edges that should be removed to simplify the network or whose edge-weight should be modified to increase or decrease the total communicability. Moreover, we would like to determine approximations of the smallest and largest elements of $L_{f}(A,ee^{T})|_{\widehat{\mathcal{L}}}$ that correspond to edges that should be added to increase total communicability. We first carry out $5$ steps of the symmetric Lanczos process, computing the matrices $\mathbf{U}_{5}$ and $\mathbf{T}_{5}$, and make use of the latter to construct both the matrix $\mathbf{X}_{5}$ in (18) by using (19), and the matrix
@@
with $h=\frac{2}{n}\cdot 10^{-4}$. Then, proceeding as described in Subsection 2.2, we found that both the matrices $\mathbf{X}_{5}$ and ${\tilde{\mathbf{X}}}_{5}$ determined the same edges. Regarding the timings (in seconds) of the two procedures, we remark that, having available the symmetric and tridiagonal matrix $\mathbf{T}_{5}$ determined by $\ell=5$ steps of the Lanczos process, evaluating the matrix $\mathbf{X}_{5}$ in (19) required the average time $t_{1}\approx 2.9\cdot 10^{-5}$ over $10^{4}$ tests, while evaluating the matrix ${\tilde{\mathbf{X}}}_{5}$ required the average time $t_{2}\approx 2.7\cdot 10^{-5}$ in the same tests, with a relative average time saving
$\frac{t_{1}-t_{2}}{t_{1}}\approx 7.0\cdot 10^{-2}.$
The smallest element of the computed approximation of $L_{f}(A,ee^{T})|_{{\mathcal{L}}}$ is $9.7462\cdot 10^{-7}$ and is associated with the edge $e(v_{123259}\leftrightarrow v_{123258})$, while the largest element is $1.5008\cdot 10^{1}$ and is associated with the edge $e(v_{19694}\leftrightarrow v_{19186})$; the smallest element of the computed approximation of $L_{f}(A,ee^{T})|_{\widehat{\mathcal{L}}}$ is $2.2047\cdot 10^{-9}$; it is associated with edge the $e(v_{25416}\leftrightarrow v_{11651})$. The largest element is $1.9380\cdot 10^{1}$ and is associated with the edge $e(v_{58080}\leftrightarrow v_{1})$.
Turning to the the structured Perron communicability, one has $P_{\mathcal{G}}^{{\mathcal{A}}}=1.9138\cdot 10^{2}$. The smallest entries of the Perron edge importance vector ${\mathcal{E}}_{\rho}$ and the relevant edges are displayed in Table 8 in the first and second columns, while the largest entries of ${\mathcal{E}}_{\rho}$ and the relevant edges are shown in the third and fourth columns. In order to increase the network communicability one should add edge $e(v_{44182}\leftrightarrow v_{44035})$, which is associated with the largest entry of the $\widehat{\mathcal{L}}$-analogue of the Wilkinson perturbation.
@@
Table 8: Example 5. The smallest entries of the Perron edge importance vector and the edges associated with the roads that could be removed in order to reduce the complexity of the network without changing the communicability in Usroads-48 significantly (displayed in the first two columns). The largest entries of the Perron edge importance vector and the edges associated with the first ten roads that should be widened/narrowed in order to increase/decrease the network communicability the most (in the third and fourth columns).
Removing the two edges in bold face in the second column of Table 8, associated with the smallest entries of the edge importance, gives the network ${\mathcal{G}}_{1}$, for which one has
$P_{{\mathcal{G}}_{1}}^{\cal A}=1.9138\cdot 10^{2};\;\;\;\;\frac{P_{\mathcal{G}}^{\cal A}-P_{{\mathcal{G}}_{1}}^{\cal A}}{P_{\mathcal{G}}^{\cal A}}=1.4851\cdot 10^{-16}.$
Setting to zero the weights associated with the two edges in bold face in the fourth column of Table 8 results in the network ${\mathcal{G}}_{2}$ with
$P_{{\mathcal{G}}_{2}}^{\cal A}=1.7487\cdot 10^{2};\;\;\;\;\frac{P_{\mathcal{G}}^{\cal A}-P_{{\mathcal{G}}_{2}}^{\cal A}}{P_{\mathcal{G}}^{\cal A}}=8.6228\cdot 10^{-2},$
while increasing the weights associated with the same edges by one results in the network ${\mathcal{G}}_{3}$, for which one has
$P_{{\mathcal{G}}_{3}}^{\cal A}=3.6445\cdot 10^{2};\;\;\;\;\frac{P^{\cal A}_{{\mathcal{G}}_{3}}-P_{\mathcal{G}}^{\cal A}}{P_{\mathcal{G}}^{\cal A}}=9.0435\cdot 10^{-1}.$
@@
$P^{\cal A}_{{\mathcal{G}}_{4}}=2.4540\cdot 10^{2};\;\;\;\;\frac{P^{\cal A}_{{\mathcal{G}}_{4}}-P^{\cal A}_{\mathcal{G}}}{P^{\cal A}_{\mathcal{G}}}=2.8228\cdot 10^{-1}.$
@@ 5 Conclusion and comments on related work
The identification of important and unimportant edges is a fundamental problem in network analysis. Several techniques for this purpose have been described in the literature; see, e.g., DLCJR ; DLCMR2 ; DLCMR3 ; EHNR ; NR ; NR2 . In DLCJNR the authors propose a method that uses the gradient of the total communicability, and Schweitzer Sc recently described how the computational effort required by this method can be reduced. Section 2 of this paper reviews this method and discusses computational aspects when this technique is applied to small and medium-sized networks, as well as to large-scale networks. In particular, further ways to speed up the computations when the method is applied to large-scale networks are described.
Another approach to identify important and unimportant edges is to determine edge weights whose modification yields a relatively large change in the Perron root of the adjacency matrix. This is described in EHNR ; NR . The computations required are quite straightforward and the method discussed in the latter reference is easy to implement also for large-scale problems. We therefore are interested in whether modifications of the weights of the edges identified by this technique give a relatively large change in the total communicability. Section 3 reviews the method described in NR and extends it to include edge removal. Computed examples reported in Section 4 show that, indeed, modifications of edge weights identified by the technique discussed in NR yield relatively large changes in the total communicability.